OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -2 and a y-intercept of 1?

OpenStudy (anonymous):

the intercept form of a line equation is \[(x/a)+(y/b)=1\]

OpenStudy (anonymous):

where a is the x- intercept and b is the y-intercept

OpenStudy (anonymous):

(x/-2)+(y/1)=1 ----->-x+2y-2=0

OpenStudy (anonymous):

I need that in i think y=ax+b form y = -x + 1 y =x + 1 y = 2x + 1